OpenStudy (anonymous):

Solve the system: 4x - y = 8 -8x + 2y = -5

OpenStudy (babalooo):

You have 2 equations and 2 unknown variables. solve one equation in terms of one variable. then plug that into the other equation. Like if y=4x-8 then plug in -8x+2 (plug in the equation here)-5 and solve for x. then once you got x, solve for the y variable using that x value

OpenStudy (anonymous):

I dont get your explanation

OpenStudy (babalooo):

what I was saying is you have 4x - y = 8 -8x + 2y = -5 then y=4x-8 -8x+2(4x-8)=-5 solve for x and once you get x, plug it back into y=4x-8 to get y